GRADUATE STUDENT COURSE ASSIGNMENTS

Block 1 Discrete Mathematics —- Babai  Weeks 1–8
Block 2a Groups and Problems —- Abert  Weeks 1–4
Block 2b Elliptic curves —- Unver  Weeks 5–6
Block 2c Complex algebraic curves —- Kerr  Weeks 7–8
Block 3a Pure and applied analysis —- Fefferman  Weeks 1–2
Block 3b Music theory —- Fiore  Weeks 3–4
Block 3c Fractals and dimension —- DeMarco  Weeks 5–6
Block 4a Problems in topology —- Farb  Weeks 1–4
Block 4b Riemann surfaces —- Souto  Weeks 5–6
Block 4c Random walks on graphs —- Bader  Weeks 7–8
Block 5 K-theory —- May  Weeks 1–2, 5–8
Block 6 Number Theory (Apprentice) —- Abert Babai Weeks 1*–4
Week 1 Week 2 Week 3 Week 4 Week 5 Week 6 Week 7 Week 8
Block 1 Block 1 Block 1 Block 1  Block 1 Block 1 Block 1
Block 2a Block 2a Block 2a Block 2a Block 2b Block 2b Block 2c Block 2c
Block 3a Block 3a Block 3b Block 3b Block 3c Block 3c
Block 4a Block 4a Block 4a Block 4a Block 4b Block 4b Block 4c Block 4c
Block 5 Block 5  Block 5 Block 5 Block 5 Block 5
*  Block 6 Block 6 Block 6

*No class meetings first week, but meetings with students

Block 1 (Babai) (weeks 1–4, 6–8) Chakraborty (CS), Turkoglu (CS); Beazley, Guichard, Schedler, Zoque-Lopez

Block 2a (Abert) (weeks 1–4) Drenning, Marcus, Noel, Thiyagarajan

Block 2bc (Unver, Kerr) (weeks 5–8) Drenning, Madapusi, Marcus, Noel, Thiyagarajan

Block 3a (Fefferman) (weeks 1–2) Carstea, Zarnescu

Block 3b (Fiore) (weeks 3–4) Biringer, Carstea

Block 3c (DeMarco) (weeks 5–6) Carstea, Dymarz, Zarnescu, Zbarsky

Block 4a (Farb) (weeks 1–4) Biringer, Booth, Diamondstone, Putman, Young

Block 4bc (Souto, Bader) (weeks 5–8) Biringer, Booth, Diamondstone, Holden, Putman

Block 5 (May) (weeks 1–2, 5–8) Abouzaid, Anders (5-8), Fowler (1-2, 5-8), Kamgarpour, Smith

Block 6 (Apprentice) (weeks 1–4, mentoring also 5–8) Anders, Booth, Holden, Madapusi, Patel, Ponto, Zbarsky

2 GRADUATE STUDENT COURSE ASSIGNMENTS

Assignments by block and by mentoring group

Name  Block Weeks Mentoring
Alan Anders  6, 5  all 8  Apprentice
Mohammed Abouzaid  5  1, 5–7 (Organizer)
Elizabeth Beazley  1  all 8  (Babai)
Ian Biringer  3b, 4  all 8  SESAME
Matthew Booth  6, 4  all 8  Apprentie
Catalin Carstea  3  all 8  YSP
David Diamondstone  4  all 8  YSP
Shawn Drenning  2  all 8  SESAME
Tullia Dymarz  3c  5–8  SESAME
Jim Fowler  6, 5 1–2, 5–8 Apprentice
Megan Guichard  1  1–4, 7–8 (Babai)
Matt Holden  6, 4bc  8  Apprentice
Masoud Kamgarpour  5  1–3, 5–8 SESAME
Keerthy Madapusi  6, 2b  1–6  Apprentice
Aaron Marcus  2  1–3, 5–8  YSP
Justin Noel  2  all 8  SESAME
Deepam Patel  6  1–4  Apprentice
Kate Ponto  6  all 8  Apprentice
Andrew Putman  4  1–6, 8 SESAME
Travis Schedler  1  all 8  (Babai)
Emma Smith  5  1–2, 5–8  YSP
Saravanan Thiyagarajan  2  all 8  SESAME
Robert Young  4a  all 8  YSP
Dani Zarnescu  3a, 3c all 8  SESAME
Boris Zbarsky  3c, 6  5–8  Apprentice
Eliana Zoque-Lopez  1  all 8  (Babai)
Sourav Chakraborty  1  all 8  (Babai)
Duru Turkoglu  1  all 8  (Babai)

ORGANIZERS

Team leaders for courses: people in charge of organizing problem sessions, making sure that graduate students are at talks, and coordinating with faculty, whose preferences on how graduate students are to help out are paramount.

1 Babai: Sourav Chakraborty

2a Abert: Justin Noel

2bc Unver-Kerr: Justin Noel 3a Fefferman: Catalin Carstea, Arghir Zarnescu 3b Fiore: Catalin Carstea 3c DeMarco: Catalin Carstea, Tullia Dymarz 4a Farb: Andy Putman 4bc Souto-Bader: Andy Putman

5 May: Masoud Kamgarpour

6 Abert-Babai (Apprentice): Kate Ponto

Mentoring organizer: in charge of setting up mentorship pairings: Abouzaid (first week: organizer in chief)

APPRENTICE: Mentors: Anders, Booth, Holden, Madapusi, Patel (weeks 1-4), Ponto, Zbarsky (weeks 5-8)

YSP: Mentors: Carstea, Diamondstone, Fowler (1-2, 5-8), Marcus, Smith (1-2, 5-8), Young

SESAME: Mentors: Biringer, Drenning, Dymarz (5-8), Kamgarpour, Noel, Putman, Thiyagarajan, Zarnescu.